OpenStudy (lilcathyx):

prove that perpendicular lines have opposite and reciprocal slopes. i just want to know how. to see if i can resolve a question...

OpenStudy (lilcathyx):

i do that by using the slope formula ? the m= \[y2-y1 \over x2-x1\]

OpenStudy (lilcathyx):

@haleyelizabeth2017 what do you think ?

OpenStudy (lilcathyx):

@phi ?

OpenStudy (phi):

I'll have to think about it. we need an idea.

OpenStudy (lilcathyx):

mhmmm.. okkay. an idea.. im sorry i dont understand this a lot.

OpenStudy (lilcathyx):

this is the question " Triangle QRS has been rotated 90° to create triangle TUV. Using the image below, prove that perpendicular lines have opposite and reciprocal slopes. You must show all of your work to receive credit. "

OpenStudy (lilcathyx):

90 degrees to create*

OpenStudy (phi):

can you post a screen shot of the picture ?

OpenStudy (lilcathyx):

Yep ;o

OpenStudy (phi):

Did you make that graph or was it given. I only ask, because the triangles are also reflected... not quite what you want.

OpenStudy (lilcathyx):

That is the graph that was given

OpenStudy (phi):

they claim the two lines SR and UV form 90 deg angle. ok, let's find the slope of line SR and the slope of UV and see if we get negative reciprocals

OpenStudy (lilcathyx):

SR = 3/3 = 1 & UV = -3/3 = -1

OpenStudy (lilcathyx):

aren't they opposite reciprocals ?

OpenStudy (lilcathyx):

i don' think they are negative reciprocals..

OpenStudy (phi):

1 is special. If we got 2 and - 1 /2 it would be obvious. but with 1, we do have negative reciprocals. But to see it, write it like this: \[ -1 = \frac{1}{-1} \] now you have \( 1 \text{ and } \frac{1}{-1} \) which are negative reciprocals. so you have shown the 90 degree lines have negative reciprocals

OpenStudy (phi):

Apparently people have introduced "opposite reciprocal" as a new way to say negative reciprocal. (who knows why ?!). But they mean the same thing: \( a \text{ and } -\frac{1}{a} \) when you multiply them you get -1.

OpenStudy (lilcathyx):

That is all i have to do to prove that perpendicular lines have opposite and reciprocal slopes ?

OpenStudy (phi):

I think that is what they want for this problem. But we have not proved that *all* perpendicular lines have opposite reciprocals, only the pair of lines in this problem. Proving it's true all the time is harder... a lot harder... so I don't think that is what they want.

OpenStudy (lilcathyx):

Oooohh i see. Thank you for helping phi :) ^.^